1. A long, straight conductor carries a current of 200 A. At what distance from the conductor is the magnetic field


caused by the current equal to the Earth’s magnetic field?



2. A power cable contains two long parallel conductors placed 1.5 cm apart. It carries a current of 125 A to a 40 hp


electric motor. Find the force exerted on a 2.0 m length of one conductor?



3. A coil used to produce a magnetic field for an electron beam experiment has a radius of 15 cm and has 300


turns. What current is needed to produce a magnetic field of 5.0 x 10-3 T?

1
Expert's answer
2022-03-21T12:46:25-0400

(1)

Magnetic field

"B=\frac{\mu_0 I}{2\pi r}"

"r=\frac{4\pi\times10^{-7}\times200}{2\pi\times 5\times10^{-5}}=0.8m"

(2) force

"F=\frac{\mu i_1i_2}{2\pi r}"


"F=\frac{4\times3.14\times 125\times125}{2\times3.14\times0.015}=2.08\times10^{6}N"

(3) magnetic field

"B=\frac{\mu_0Ni}{2r}"

"i=\frac{5\times10^{-3}\times2\times0.15}{300\times4\times3.14\times10^{-7}}=3.98A"